how do i make this mixed fraction 14 5/12 a decimal??
Answer by ikdeep(226)

first convert this mixed fraction into proper fraction anad its done like this ,....
numerator = 12* 14 + 5
= 168 + 5
= 173

denomiator is as such = 12
and u get proper fraction = 173/12
now divide 173 by 12 and you will get answer in decimal form..
Hope this will help you
Please feel free to revert back for any further queries.